OpenStudy (anonymous):

I know the answer is 3 I just can't figure out how they got it :p

OpenStudy (anonymous):

\[9\sqrt{13a-3}=54\]

OpenStudy (amistre64):

divide, square, add, then divide again

OpenStudy (anonymous):

Divide what?

OpenStudy (amistre64):

these are basic and fundatmental algebraic techniques ... 9x = 54, what would you divide to find x?

OpenStudy (amistre64):

...

OpenStudy (anonymous):

You'd divide 54 by 9 in order to find x

OpenStudy (amistre64):

good, then x = 54/9 = 6 move to the next "portion" sqrt(x) = 6 how would we determine an x for this?

OpenStudy (anonymous):

There's no x in the problem? nd the variable is under the square root symbol?

OpenStudy (amistre64):

im taking you thru the steps that you would do if you had seperate problems

OpenStudy (amistre64):

how would you solve sqrt(x) = 6 ?

OpenStudy (anonymous):

Idk

OpenStudy (amistre64):

divide, square, add, then divide again we already divided off the 9, the next process is to ?

OpenStudy (anonymous):

Take the square root of x?

OpenStudy (amistre64):

almost, take the sqaure of sqrt(x) and the square of 6 x = 6^2 = 36

OpenStudy (amistre64):

now we are left with: 13x - 3 = 36 how would we complete this process?

OpenStudy (anonymous):

wolf will you finish typing already? nd Then you add?

OpenStudy (amistre64):

9 sqrt(13-3) = 54 \9 \9 ----------------- sqrt(13a - 3) = 6 , square each side 13a - 3 = 36 etc ...

OpenStudy (amistre64):

yes, what do we add?